I. Two cards are drawn from a well-shuffled ordinary deck of 52 cards. Find the probability that they are both aces if the first card is  a) replaced,  b) not replaced. II. A ball is drawn at random from a box containing 6 red balls, 4 white balls, and 5 blue balls. Determine the probability that it is (a) red, (b) white, (c) blue, (d) not red, (e) red or white. III. Prove that the mean and variance of a binomially distributed random variable are, respectively, µ = np and σ2 = npq. 


1
Expert's answer
2021-06-17T17:57:45-0400

(1). Probability of drawing two aces from 52 cards is-


(a) If first is replaced-


"P=\dfrac{4}{52}\times \dfrac{4}{52}=\dfrac{1}{169}"


(b) If first is not replaced-


"p=\dfrac{4}{52}\times \dfrac{3}{51}=\dfrac{1}{221}"


(2). Given, 6 Red balls, 4 White balls, 5 blue balls-

Total balls=15


"(a) P(\text{Red ball})=\dfrac{6}{15}=\dfrac{2}{5}


\\[9pt]
 (b) P(\text{White ball})=\dfrac{4}{15}\\[9pt]



 (c) P(\text{Blue ball})=\dfrac{5}{15}=\dfrac{1}{3}\\[9pt]



 (d) P(\text{not a Red ball})=1-P(Red ball)=1-\dfrac{2}{5}=\dfrac{3}{5}
\\[9pt]


 (e) P( \text{Red or White})= P(R)+P(W)=\dfrac{6}{15}+\dfrac{4}{15}=\dfrac{10}{15}=\dfrac{2}{3}"


(3).

Mean="Ex=\sum x \text{ }^nC_x p^xq^{(1-x)}=np"

"x\times \dfrac{n!}{(n-x)! x!}=x\dfrac{n(n-1)!}{(n-x)! x (x-1)!}=n\text{ }^{n-1} C_{x-1}"


So Mean "= \sum n\text{ }^{n-1}C_{x-1}p^{x-1}.p.(1-p)^{n-x}"


   "np\sum ^{n-1}C_{x-1}"


Let x-1=y, n-1=m


So Mean="np \sum ^mC_y P^y(1-p)^{m-y}"


As we know "\sum f(x)=1"


So Mean"=np\times 1=np"



Variance "= E[x^2]-[E(x)]^2=\sum x(x-1) \text{ }^n C_xp^xq^{(1-x)}"



      

So "E[x(x-1)]=\sum x(x-1) \text{ } ^nC_xp^xq^{n-x}"


       


  "x(x-1) \text{ } ^nC_x=x(x-1)\dfrac{n!}{(n-x)! x!}=x(x-1)\dfrac{n(n-1)(n-2)!}{(n-x)!x(x-1)(x-2)!}"

       

"=n(n-1)\dfrac{(n-2)!}{(n-x)!(x-2)!}

\\[9pt]

 =n(n-1)\text{ } ^{n-2}C_{x-2}"


So "E[x(x-1)]=\sum n(n-1) \text{ } ^{n-2}C_{x-2} p^{x-2} (1-p)^{n-x}"

      

"=(n(n-1))p^2 \sum \text{ } ^{n-2}C_{x-2} p^{x-2}(1-p)^{n-x}"

           Let x-2=y,n-2=m


     "=n(n-1)p^2 \sum \text{ } ^mC_y p^y(1-p)^{m-y}"


      "=n(1-n)p^2 \times 1"   


As "\sum f(x)=1"


Variance="E(X^2)-[E(x)]^2"

     "=E[x(x-1)]+E(X)-[E(x)]^2
\\
 =n(n-1)p^2+np-(np)^2
\\
 =n^2p^2-np^2+np-n^2p^2
\\
 =np(1-p)"